# Modelling Coincident Particle Production in Ultraperipheral Heavy Ion Collisions

### v1

@@ 4 Conclusions
In this paper we have considered the case of coincident particle production in ultraperipheral heavy ion collisions, that is the production of some state of interest in addition to the primary UPC process under study, due to multiple initial–state photon emission. We have focussed on the case of photon–initiated dimuon production with coincident $\rho^{0}$ meson photoproduction in PbPb collisions, which is included in the SuperChic MC generator, and will be publicly available at
@@
This is motivated by the recent ATLAS measurement [8] of this process. We have presented the first precise theoretical calculation of the expected rate for this, and the non–trivial dependence of it on the ZDC selection of the colliding ions as well as the kinematics of the primary dimuon state. Results for this are found to be in very encouraging agreement with the data, up to an overall normalization of the $\gamma A\to\rho^{0}A$ subprocess cross section, which is in some tension with the expectations based on ALICE data of primary $\rho^{0}$ production in UPCs. The non–trivial dependence of the coincidence rate on the ZDC selection and dimuon kinematics is in particular reproduced well.
@@
We have in addition considered the possible impact of such coincident production more generally on standard UPC measurements. In particular, as an exclusivity veto is applied in this cases, any such coincident production that results in additional particles in the veto region will be rejected. As this is not currently corrected for, and is not included in existing theoretical calculations, this will result in some overshoot of the theory in comparison to data. As discussed in this paper, due to the kinematic and ZDC class dependence of the coincidence probability, this effect will not be constant but will rather depend on these variables.
@@ Acknowledgements
I thank Brian Cole for useful discussions on this topic, and Soumya Mohapatra for providing the data corresponding to the ATLAS measurement. I thank the Science and Technology Facilities Council (STFC) part of U.K. Research and Innovation for support via the grant award ST/T000856/1.

### v2

@@ 4 Conclusions
In this paper we have considered the case of coincident particle production in ultraperipheral heavy ion collisions, that is the production of some state of interest in addition to the primary UPC process under study, due to multiple initial–state photon emission. We have focussed on the case of photon–initiated dimuon production with coincident $\rho^{0}$ meson photoproduction in PbPb collisions, which is included in the SuperChic MC generator, and is publicly available at
@@
This is motivated by the recent ATLAS measurement [8] of this process. We have presented the first precise theoretical calculation of the expected rate for this, and the non–trivial dependence of it on the ZDC selection of the colliding ions as well as the kinematics of the primary dimuon state. Results for this are found to be in very encouraging agreement with the data. The non–trivial dependence of the coincidence rate on the ZDC selection and dimuon kinematics is in particular reproduced well.
@@
We have also considered the possible impact of such coincident production more generally on standard UPC measurements. In particular, as an exclusivity veto is applied in this cases, any such coincident production that results in additional particles in the veto region will be rejected. If this is not currently corrected for, then given it is not included in existing theoretical calculations, this will result in some overshoot of the theory in comparison to data. As discussed in this paper, due to the kinematic and ZDC class dependence of the coincidence probability, this effect will not be constant but will rather depend on these variables.
@@ Acknowledgements
I thank Valery Khoze and Misha Ryskin for reading the manuscript and providing comments. I thank Daniel Tapia Takaki for bringing to my attention the issue of the non–resonant dipion contribution to the ATLAS data. I thank Brian Cole and Ronan McNulty for useful discussions on this topic, and Soumya Mohapatra for providing the data corresponding to the ATLAS measurement. I thank the Science and Technology Facilities Council (STFC) part of U.K. Research and Innovation for support via the grant award ST/T000856/1.